A culture of bacteria has 20 cells in it. The number of cells doubles every 30 minutes
A) write the first 5 terms of this sequence
B)is this sequence arithmetic or geometric? Explain how you know
C)write the general term for the sequence
D)use your formula to determine the number of cells after 8 hours
Step by step answers would be appreciated thanks in advance ☺

Answer by Alan3354(69443)

A culture of bacteria has 20 cells in it. The number of cells doubles every 30 minutes
-------------
A) write the first 5 terms of this sequence
Using 30 mins as the period:
20, 40, 80, 160, 320
------------------
B)is this sequence arithmetic or geometric? Explain how you know
Geometric. The terms are multiplied, not added to.
------------------------
C)write the general term for the sequence
C = 20*2^n where C = # of cells, n = # of periods
--------------
D)use your formula to determine the number of cells after 8 hours
8 hours = 16 periods
C = 20*2^16
C = 1310720 cells